OpenStudy (anonymous):

(-2a^3b^2)(ab^(-4)/a^(-3))^3

OpenStudy (anonymous):

\[\large\left( -2a^3b^2 \right)\left( \frac{ ab^\left( -4 \right) }{ a^\left( -3 \right) } \right)^3\]